FM SPECTROSCOPY USING PULSE-AMPLIFIED LASERS

\underline{E. E. EYLER} and A. N. NIKOLOV, Physics Department, University of Connecticut, Storrs, CT 06269.

We have explored techniques for sensitive FM absorption spectroscopy with nanosecond laser pulses, both in a simple absorption cell and in a resonant enhancement cavity. In both cases, intense transform-limited pulses of frequencymodulated light are produced by pulsed amplification of an FM modulated cw laser. In our first experiments, performed in collaboration with Robert Field and Jonathan Bloch from MIT, ^{a} we measured FM absorption in the γ band of NO. To produce the required 214 nm radiation, pulsed FM light was frequency-tripled using KDP and BBO crystals. A fractional sensitivity of 10^{−4} was achieved in single-pass absorption. More recently we have extended the pulsed FM technique by adding a resonant cavity. Transform limited pulses of frequency modulated light are injected into a locked, single mode cavity of moderate finesse, matched to the sideband spacing. Preliminary results on iodine demonstrate an absorption sensitivity of roughly 1 × 10^{−9} per cm, showing promise for future applications in far UV spectroscopy, Doppler-free saturation spectroscopy, and multiphoton absorption spectroscopy.

aE.E. Eyler, S. Gangopadhyay, N. Melikechi, J.C. Bloch, and R.W. Field, Opt. Lett. 21, 225 (1996)